MTA Monitoring Report 09/16/05 - 09/22/05

Archive of all weekly reports

Archive of monthly reports

At the 259:0345 pass the OC/CC noticed that the DEA-A was off. A quick analysis shows that the DEA-A voltages and currents were nominal up until the major frame at which the DEA shut off. The ACIS team assessment is that this looks like the two cases where DPA turned off. Those were not assigned a definitive cause, but the best hypothesis was a cosmic ray hit to the opto-isolator causing the PSMC to send a spurious command to turn off.



Detrended CTI

CTI page was updated. The top panel shows an example of detrended CTI measurements from the front illuminated CCD3 Node 0. The top plot is all data points without any temperature dependency correction. The middle plot is data with the focal temperature equal or colder than -119.7 C, and the integration time longer than 7,000 sec. The bottom plot is all data with the temperature correction.

The values quoted below are temperature factor corrected ones.
Average of ACIS-I CCDs (MnKa) slope 4.378e-9 CTI/day
Detrended slope 7.696e-9 CTI/day

ACIS Warm Pixels

We display all bad pixels showed up in past, and, if any, previously unknown bad pixels appeared in the last 14 days. We also list hot pixels (defined as 1000 above the average of bias background). Warm columns 509-514 are probably due to a computational artificial effect due to boundaries.

ACIS Focal Plane Temperature

SIM Movements

9 TSC moves this period
weekly average time/step 0.00149 s
mission average time/step 0.00136 s



Recent Observations

Temporarily unavailable.


This Week's focus is HRC.

Trends last reported on June 30.

Only the most interesting or representative msids are shown below. For a full listing choose the bulletted link.